OpenStudy (halfdeafdarling):

how do i find the restrictions of a simplified expression? like for 4x^2-25/2x-5 what is the restriction? I know the answer which when solved is 2x+5 but where would i begin to find out the restriction?

jimthompson5910 (jim_thompson5910):

Rule: you cannot divide by zero something like 1/0 is undefined

jimthompson5910 (jim_thompson5910):

since 0 cannot be in the denominator, this means that 2x-5 cannot be zero if it were zero, then 2x-5 = 0 leads to x = ???

OpenStudy (halfdeafdarling):

2x=5

OpenStudy (halfdeafdarling):

/2 ohhhh

OpenStudy (halfdeafdarling):

THANK YOU

jimthompson5910 (jim_thompson5910):

5/2 or 2.5 if x = 5/2 or x = 2.5, then the denominator becomes zero. So that's why you have to place the restriction

OpenStudy (halfdeafdarling):

oh okay thank you so much that explains it much better!

jimthompson5910 (jim_thompson5910):

you're welcome